OpenStudy (anonymous):

if the linear equation 4x=10 -5y is written in standard form, what is the value of B?

OpenStudy (anonymous):

please help :(

OpenStudy (anonymous):

Standard form = Ax + By = C

OpenStudy (anonymous):

do you know how to change the formula into standard form?

OpenStudy (anonymous):

not a clue!!

OpenStudy (anonymous):

Well slope intercept form is y = mx + b

OpenStudy (anonymous):

so your formula now is : 4x = 10 - 5y you just need to get 5y on the other side. :P

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

your equation should look like 4x + 5y = 10

OpenStudy (anonymous):

B = 5

OpenStudy (anonymous):